Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280 lb mixture that is 12% protein? : Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280 lb mixture that is 12% protein?
Answer by checkley77(12569)

.14x+.07(280-x)=.12*280
.14x+19.6-.07x=33.6
.07x=33.6-19.6
.07x=14
x=14/.07
x=200 lb. of soybeans are used.
280-200=80 lb. of cornmeal is used.
Proof:
.14*200+.07*80=33.6
28+5.6=33.6
33.6=33.6